OpenStudy (anonymous):

Will Someone Please Help Me//: If so thanks <3 (ill Fan) The table below shows two equations: Equation 1 |4x - 3|- 5 = 4 Equation 2 |2x + 3| + 8 = 3 Which statement is true about the solution to the two equations?

OpenStudy (anonymous):

Equation 1 and equation 2 have no solutions. Equation 1 has no solution and equation 2 has solutions x = -4, 1. The solutions to equation 1 are x = 3, -1.5 and equation 2 has no solution. The solutions to equation 1 are x = 3, -1.5 and equation 2 has solutions x = -4, 1.

OpenStudy (anonymous):

@jigglypuff314 you think you could help me?

OpenStudy (ranga):

|4x - 3| - 5 = 4 add 5 to both sides: |4x - 3| = 9 which implies: 4x - 3 = 9 or 4x - 3 = -9 4x-3 = 9 4x = 12 x = 3 4x-3 = -9 4x = -6 x = -6/4 = -3/2 Therefore, the solution to equation 1 are: 3, -3/2 Equation 2:

OpenStudy (ranga):

2 |2x + 3| + 8 = 3 add -8 to both sides: 2|2x+3| = -5 divide both sides by 2 |2x+3| = -5/2 But left hand side is always positive because of the absolute bars but thee right hand side is negative. Therefore, Equation 2 has no solution.

OpenStudy (anonymous):

thank you

OpenStudy (ranga):

you are welcome.